# jmstate, a Flexible Python Package for Multi-State Joint Modeling

### v1

@@ 5 Likelihood and Statistical Inference
5.1 Stochastic Gradient Estimation via Fisher Identity
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [Kuhn and Lavielle, 2004], Laplace approximation [Wolfinger, 1993], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [Robbins and Monro, 1951], and MCMC-based approximations such as Metropolis–Hastings and Hamiltonian Monte Carlo. These strategies are implemented in software such as JMBayes [Rizopoulos et al., 2024, Rizopoulos, 2020]. Another approach requiring mild regularity assumptions on the log marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [Caillebotte et al., 2025, Baey et al., 2023] using Fisher’s identity and following the Robbins-Monro procedure [Robbins and Monro, 1951].
Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes:
$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{full}}(\theta;x,b)\right).$
@@ 5.1 Stochastic Gradient Estimation via Fisher Identity
1:data $x$; initial parameter $\theta_{0}$; step sizes $(\eta_{n})_{n\geq 0}$ with $\sum_{n}\eta_{n}=+\infty$, $\sum_{n}\eta_{n}^{2}<+\infty$; batch size $K$; sampler $p_{n}^{\mathrm{MCMC}}$; stopping criterion.
@@
5:  $\theta_{n+1}\leftarrow\theta_{n}+\frac{\eta_{n}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}(\theta_{n};x,b_{k})$.
@@
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [Kingma and Ba, 2014], NAdam [Dozat, 2016] or Adagrad [Duchi et al., 2011]. In particular, the Adam optimizer will be later used in our numerical experiments.
6 Implementation Details
The nonlinear joint model is fully implemented in Python, with PyTorch and NumPy as its only major dependencies. Notably, the stochastic gradient scheme is carried out following Algorithm 2 and heavily relies on PyTorch’s automatic differentiation and broadcasting capabilities [Paszke et al., 2017].
6.0.0.1 Posterior Sampling and Parameter Estimation.
A Metropolis-Hastings sampler is used to simultaneously draw posterior samples of the random effects in an efficient way. Adaptive step sizes are implemented at the individual level via the Robbins-Monro procedure. We favored Metropolis-Hastings over alternatives such as Hamiltonian Monte Carlo or Metropolis-within-Gibbs for its simplicity. Autocorrelation can be monitored and mitigated by subsampling. The acceptance rate is targeted at $0.234$, which is asymptotically optimal for Gaussian proposals in high dimensions [Gelman et al., 1997]. Although our posteriors are not Gaussian, this rule is simple and effective. Moreover, as the number of longitudinal measurements increases ($m_{i}\to+\infty$), posterior distributions converge to a multivariate normal (Rizopoulos et al. [2008]).
6.0.0.2 Numerical integration.
Cumulative hazards are computed using Gauss–Legendre quadrature [Lether, 1978], which integrates exactly all polynomials up to degree 2n12n-1 with $n$ nodes:
$\int_{a}^{b}f(t)\,dt\approx\sum_{i=1}^{n}f(t_{i})w_{i}.$
We set $n=32$ by default. Since nodes are fixed, integration points are cached, using a hash table, outside the backpropagation graph, reducing cost and memory. The positivity of the quadrature weights ensures stable computation and preserves the positivity of integrals.
6.0.0.3 Simulation of event times.
Event times are generated by inverting the survival function via a bisection search. An exponential random variable is first drawn to determine the event threshold. Starting from the initial and censoring times, the cumulative hazard is iteratively approximated with Gauss–Legendre quadrature until the threshold is reached. This procedure naturally incorporates censoring and yields the simulated event time.
6.0.0.4 Covariance matrix parameterization.
@@ 7 Simulation Study
To study the robustness of the estimator, we propose a simulation example and compute an estimate of the Root Mean Square Error (RMSE) for each parameter, based on different datasets generated from a common underlying distribution.
In parallel, we illustrate the usage of the jmstate package through concise code snippets, structured as follows:
  • Specification of the true underlying model.
  • Simulation of longitudinal and event data.
  • Stochastic optimization of the parameters with corresponding diagnostic plots.
@@ 7.1 Model Specification
The model considered involves three states (See Figure 3), labeled 0 (healthy), $1$ (sick), and $2$ (terminal stage). The allowed transitions are $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$, corresponding to a biological modeling of the biomarker evolution, where a too rapid decrease of the biomarker or a too low value leads to the degradation of the patient’s health status.
Figure 3: State transition diagram of the multi-state model: transitions are allowed from Healthy (0) to Sick (11), from Healthy (0) to Terminal Stage (22), and from Sick (11) to Terminal Stage (22).
The regression function is a piecewise affine approximation of a bi-exponential curve, with an inflexion point at a fixed known time $\tau=6$ which may correspond to a drug administration time, causing a rupture in the longitudinal behavior while remaining continuous:
@@
which can be written in Python, after necessary imports, as:
import torch TAU = 6.0 def reg(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) # Extract relevant terms # psi has shape (n_chains, n_individuals, n_repetitions) # output has shape (n_chains, n_individuals, n_repetitions, 1) return (b + w1 * t + (w2 - w1) * (t > TAU) * (t - TAU)).unsqueeze(-1)
Note this python function must accept tensors psi with two or three dimensions and return a tensor of dimension $\operatorname{dim}(\text{psi})+1$, with a shape whose last dimension matches the dimension of the biomarker (here, $\mathbb{R}^{1}$ so it is equal to $1$).
For the link function, the concatenation of $h$ as well as its partial derivative with respect to time is considered and shared across all possible transitions: $g=\left(h,\frac{\partial h}{\partial t}\right)$, which is efficiently implemented:
def link(t: torch.Tensor, psi: torch.Tensor): b, w1, w2 = psi.chunk(3, dim=-1) diff = (w2 - w1) * (t > TAU) val = b + w1 * t + diff * (t - TAU) der = w1 + diff return torch.cat([val.unsqueeze(-1), der.unsqueeze(-1)], dim=-1)
Additionally, normally distributed covariates $X$ and exponential baseline hazards were considered with a clock reset specification. The model design is therefore declared as follows:
from jmstate.functions import Exponential, gamma_plus_b from jmstate.typedefs import ModelDesign # Survival model specification surv = { (0, 1): (Exponential(0.1), link), (0, 2): (Exponential(0.01), link), (1, 2): (Exponential(0.2), link), } # Model design gathers regression, link and hazard functions model_design = ModelDesign(gamma_plus_b, reg, surv)
@@
The model is therefore parameterized in $\mathbb{R}^{16}$ and the true parameters may be defined as follows:
from jmstate.typedefs import ModelParams from jmstate.utils import repr_from_cov # Gaussian means gamma = torch.tensor([2.5, -1.3, 0.2]) # Covariance matrices Q = torch.diag(torch.tensor([0.6, 0.2, 0.3])) R = torch.tensor([[1.7]]) # Link parameters alphas = { (0, 1): torch.tensor([-0.5, -3.0]), (0, 2): torch.tensor([-1.0, -5.0]), (1, 2): torch.tensor([0.0, -1.2]), } # Covariate parameters betas = { (0, 1): torch.tensor([-1.3]), (0, 2): torch.tensor([-0.9]), (1, 2): torch.tensor([-0.7]), } # Instance declaration real_params = ModelParams( gamma, repr_from_cov(Q, method="diag"), repr_from_cov(R, method="ball"), alphas, betas, )
Note the argument method, which can take values "full", "diag", or "ball". It represents the chosen parameterisation of the covariance matrices, either unconstrained (full), diagonal (diag) or scalar (ball).
7.2 Data Generation
With the parameter values defined above, as well as the choice of the regression and link functions, we are now able to declare the true base model and use it to simulate longitudinal processes and transition trajectories, with the help of the following helper function:
from jmstate import MultiStateJointModel from jmstate.typedefs import ModelData, SampleData from torch.distributions import MultivariateNormal # Declares the true underlying model real_model = MultiStateJointModel(model_design, real_params) def gen_data(n: int, m: int): # Censoring times c = torch.rand(n, 1) * 5 + 10 # Covariates x = torch.randn(n, 1) # Latent and noise distributions Q_dist = MultivariateNormal(torch.zeros(Q.size(0)), Q) R_dist = MultivariateNormal(torch.zeros(R.size(0)), R) # Individual effects b = Q_dist.sample((n,)) psi = model_design.individual_effects_fn(gamma, x, b) # Generates random evaluations points with a minimum distance a = torch.zeros((n, 1)) b = torch.full((n, 1), 15) t = random_far_apart(n, m, a, b, 0.7 * b / m) # Defines initial state for individuals trajectories_init = [[(0.0, 0)]] * n # Samples trajectories sample_data = SampleData(x, trajectories_init, psi) trajectories = real_model.sample_trajectories(sample_data, c) # Samples longitudinal values y = model_design.regression_fn(t, psi) y += R_dist.sample(y.shape[:2]) # Censors longitudinal measurements exceeding censoring times y[t > c] = torch.nan return x, t, y, trajectories, c # Generate data data = ModelData(*gen_data(1000, 20))
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below. The function build_buckets was used to determine the number of transitions observed in the dataset. Given a list of trajectories, it returns a dictionary of transitions as keys and Buckets(NamedTuple) as values, containing the individual indices and times of the transitions.
@@ 7.2 Data Generation
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the full population of $1000$ individuals.
@@ 7.3 Estimator Convergence
To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved. In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices:
# Declares initial parameters; zero mean and unit variance init_params = ModelParams( torch.zeros_like(gamma), repr_from_cov(torch.eye(Q.size(0)), method="diag"), repr_from_cov(torch.eye(R.size(0)), method="ball"), {k: torch.zeros_like(v) for k, v in alphas.items()}, {k: torch.zeros_like(v) for k, v in betas.items()}, )
To run the optimization process, we can make use of the Fit class that provides a simple interface for choosing the optimizer (torch.optim.Adam by default) and set its hyperparameters through the kwargs argument. Additionally, a stochastic stopping condition based on the first and second moments of the parameters’ consecutive differences can be easily employed with the StopParams class.
@@
$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}\coloneqq\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
and the optimization process is terminated when:
@@
At last, the LogParamsHistory class can be used to log the parameters’ history, allowing to monitor the convergence of the parameters.
This is all done automatically by executing the do method of the model which runs the Markov Chain Monte Carlo (MCMC) algorithm to sample from the posterior distribution of the parameters based on the inputed data, a list of Job factories (such as fitting, logging, stopping, or even prediction once the model is fitted). It optionnaly returns metrics collected by the Job factories. The hyperparameters such as the target acceptance rate, the number of chains, the warmup period or the subsample parameter are automatically set based on the job_factories provided but can be overriden:
from jmstate.jobs import Fit, LogParamsHistory, ParamStop # Declares initial model model = MultiStateJointModel(model_design, init_params) # Runs optimization process metrics = model.do( # Metrics object data, job_factories=[ Fit(lr=0.5, fused=True), ParamStop(rtol=0.1), LogParamsHistory(), # Returns a list of ModelParams ], max_iterations=500, )
Although not demonstrated here, the model also supports gradient clipping, scheduling, stacking multiple other stopping criteria based on likelihood or gradient magnitude, penalized likelihood, as well as covariate selection. For more details, please refer to the official documentation of the jmstate package.
The following figure shows the evolution of the parameters during the optimization process and Table 1 summarizes the accuracy of the inferred parameters.
(a) $\gamma$
(b) $Q$
(c) $R$
(d) $\alpha^{1\mid 0}$
(e) $\alpha^{2\mid 0}$
(f) $\alpha^{2\mid 1}$
(g) $\beta^{1\mid 0}$
(h) $\beta^{2\mid 0}$
(i) $\beta^{2\mid 1}$
@@
Table 1: Comparison of true and inferred parameters (up to 4 decimal places). RMSE is computed using the bias-variance identity.
The algorithm demonstrates good adaptability and robustness, with a notably low bias. The fitting is also very quick to perform, in about $10$ to $15$ seconds per run on an AMD Ryzen 5 5600 processor, despite the use of mini-batch optimization and the higher complexity of the model induced by the nonlinearity of its functions as well as the presence of numerous possible transitions. It is also to be noted that a minimum number of observations for each transition is required to ensure the stability of the optimization process.
Finally, the convergence patterns shown in Figure 5 are remarkably smooth, a result due to the parallelization of multiple chains and the use of mini-batch optimization.

### v2

@@ 4 Statistical Inference
Several optimization methods adapted from the nonlinear mixed-effects literature can be applied, including Stochastic EM [kuhn2004coupling], Laplace approximation [wolfinger1993laplace], Gauss-Hermite quadrature, stochastic gradient ascent with Robbins–Monro updates [robbins1951stochastic], and MCMC-based approximations such as Metropolis-Hastings and Hamiltonian Monte Carlo. These strategies are implemented in software such as JMBayes [jmbayes22024, rizopoulos2020package].
Another approach requiring mild regularity assumptions on the $\log$ marginal likelihood, but without the need for the model to belong in the exponential family, is to consider a stochastic gradient ascent scheme [caillebotte2025estimation, baey2023efficient] using Fisher’s identity and following the Robbins-Monro procedure [robbins1951stochastic].
Under interchangeability of integral and differentiation, setting $x=(X,Y,\mathcal{T},C)$ for convenience, the Fisher identity writes
$\nabla_{\theta}\log\mathcal{L}_{\text{marginal}}(\theta;x)=\mathbb{E}_{b\sim p(\cdot\mid x,\theta)}\left(\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta;x,b)\right).$
@@
1:data $x$; initial parameter $\theta^{(0)}$; step sizes $(\eta_{t})_{t\geq 0}$ with $\sum_{t}\eta_{t}=+\infty$, $\sum_{t}\eta_{t}^{2}<+\infty$; batch size $K$; sampler $p_{t}^{\mathrm{MCMC}}$; stopping criterion.
@@
5:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\frac{\eta_{t}}{K}\sum_{k=1}^{K}\nabla_{\theta}\log\mathcal{L}_{\text{complete}}(\theta^{(t)};x,b_{k})$.
@@
7:return $\hat{\theta}$
Preconditioning matrices may also be used, such as the Fisher Information Matrix in a natural gradient descent framework, or even specialized optimizers such as Adam [kingma2014adam], NAdam [dozat2016incorporating] or Adagrad [duchi2011adaptive]. In particular, the Adam optimizer will later be used in our numerical experiments.
5 Dynamic prediction
The simulation of individual trajectories using Algorithm 2 enables prediction in the multi-state joint modeling framework. Specifically, predictions are derived from simulated event trajectories, allowing for individualized risk assessment and forecasting of future states or event times based on a subject’s observed biomarker and event history. This approach generalizes classical dynamic prediction by leveraging the rich structure of multi-state processes and the joint distribution of longitudinal and event data.
Suppose we are interested in some functional of the true (unobserved) future trajectory,
$\chi^{*}(\mathcal{T}_{i}^{*})\in\mathbb{Y},$
where $\mathbb{Y}$ is some generic space.
For example, one can think of the state taken by the individual $i$ at a time $u$. As is the case in traditional dynamic prediction however, we are only interested in predicting quantities or characteristics that depend on a yet unobserved future, given prior information, i.e. the trajectory up to some prediction time $t\leq u$ and observed longitudinal markers. Contrary to (single transition) joint models, the conditional probability distribution on $(\mathbb{R}\times\mathcal{S})^{\mathbb{N}}$ cannot be analytically derived.
Nonetheless, as seen in Algorithm 2, we are able to accurately simulate each trajectory. Therefore, a Monte Carlo estimation scheme may be devised under certain restrictions, where we estimate $\chi^{*}(\mathcal{T}_{i}^{*})$ by its expected value:
$\hat{\chi}_{i}=\mathbb{E}_{\mathcal{T}^{\prime}_{i}\sim\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t)}\left(\chi^{*}(\mathcal{T}^{\prime}_{i})\right)\approx\frac{1}{B}\sum_{k=1}^{B}\chi^{*}(\mathcal{T}_{i}^{(k)}),\,\mathcal{T}_{i}^{(k)}\overset{\text{i.i.d.}}{\sim}\mathscr{L}(\mathcal{T}_{i}^{*}\mid\mathcal{Y}(t),\mathcal{T}_{i},t).$
In general however, $\chi^{*}$ may very well depend on (countably) infinitely many transitions. As a result, since the proposed estiamtion relies on simulted samples, we require the quantity to depend only on a finite subset of transitions.
Assumption 1.
Assume there exists $\chi:\bigcup_{n\geq 1}(\mathbb{R}\times\mathbb{S})^{n}\to\mathbb{Y}$ and $\tau_{i}$ a stopping time for the filtration $\mathcal{F}_{in}=\sigma\left((T_{il},S_{il})_{0\leq l\leq n}\right)$ such that:
$\begin{cases}\tau_{i}<+\infty\text{ a.s.},\\ \chi^{*}(\mathcal{T}_{i}^{*})=\chi\left((T_{il},S_{il})_{1\leq l\leq\tau_{i}}\right)\end{cases}$
Essentially, given Assumption 1, with probability one we are guaranteed that the quantity of interested may be computed in a finite number of simulation steps. Indeed, the prediction algorithm for a single sample may be summarized as follows:
@@ 5 Dynamic prediction
2:
[…]
4:while $n<\tau_{i}$ do
5:  Append simulated $(T_{in},S_{in})$ to $\mathcal{T}_{i}$ if $n>\operatorname{len}(\mathcal{T}_{i})$
6:  $n\leftarrow n+1$
7:end while
8:return $\chi(\mathcal{T}_{i})$
Numerous quantities of interest may be encompassed by this framework. We give multiple examples below. First, for a directed acyclic graph $G=(V,E)$, we define the depth of $G$, denoted by $\operatorname{depth}(G)$, as the length of the longest directed path in $G$, that is,
$\operatorname{depth}(G)=\max_{(v_{0},v_{1},\dots,v_{k})}k,$
where $(v_{0},v_{1},\dots,v_{k})$ ranges over all directed paths in $G$.
Example 1 (State at time $u$).
Let $G=(V,E)$ be a directed acyclic graph, $u\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq u\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. Then, take $\chi_{u}^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $u$. In particular, if $V=\{0,1\}$ and $E=\{(0,1)\}$, we recover the special case used for survival probability estimation in standard joint models [rizopoulos2011dynamic].
Example 2 (Hitting time).
Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\vee\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$. Then, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$ and take $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})=T_{i\tau_{i}}+\mathds{1}_{\{S_{in}\}\not\rightsquigarrow\mathbb{A}}(+\infty)$ represents the hitting time for the set $\mathbb{A}$.
In Example  1, the stopping time $\tau_{i}$ captures the step at which the process for individual $i$ reaches or exceeds a given time $u$, or enters an absorbing state. The resulting value $\chi^{*}(\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$ therefore represents the state of the individual at that specific time.
In Example 2, the stopping time $\tau_{i}$ corresponds to the first time the process reaches a target subset of states $\mathbb{A}$, or becomes unable to reach it in the future. The associated value $\chi_{\mathbb{A}}^{*}(\mathcal{T}_{i}^{*})$ thus represents the hitting time of the set $\mathbb{A}$, which may be finite if $\mathbb{A}$ is reached, or $\infty$ otherwise.
Other practical applications may include finding the expected number of edges in some given trajectory, the number of times an individual has returned in a specific state, the expected time between transitions…
This simulation-based dynamic prediction framework generalizes classical approaches from single-event joint models to the multi-state context. It enables individualized forecasting of future state occupancy, event risks, sojourn times, and other clinically relevant outcomes, fully exploiting the subject’s observed biomarker trajectory and event history. In Section 7, we demonstrate its application to forecasting dependency trajectories in the PAQUID cohort.
@@ 6 Simulation Study
The purpose of this simulation study is to evaluate the finite-sample performance, identifiability, and convergence properties of the proposed inference algorithm under controlled conditions where the true parameters are known. We simulate data from a three-state semi-Markov process coupled with a nonlinear longitudinal biomarker trajectory, representing a simplified disease progression model. This setup allows us to assess both statistical accuracy and computational scalability of the stochastic-gradient estimation procedure.
All simulations, estimations, and figures reported in this section were produced using the open-source jmstate Python package, available on PyPI, which we developed to implement the multi-state joint modeling framework and inference algorithms described in Sections 3.2 and 4.
@@ 6.1 Model Specification
The simulated model involves three states—Healthy (0), Sick ($1$), and Terminal (22)—connected by the transitions $0\rightarrow 1$, $0\rightarrow 2$, and $1\rightarrow 2$ (Figure 3). This configuration captures the typical monotone evolution of a disease process, with both direct and indirect paths to the terminal state. Transition intensities follow an exponential baseline hazard under a clock-reset convention.
The longitudinal process follows a continuous piecewise-affine regression:
@@
with a known inflection time $\tau=6$. The individual-specific parameters $\psi_{i}=(\psi_{i1},\psi_{i2},\psi_{i3})$ arise from a nonlinear mixed-effects model with Gaussian random effects $b_{i}\sim\mathcal{N}(0,Q)$ and independent Gaussian noise $\epsilon_{ij}\sim\mathcal{N}(0,R)$. The link function combines the current biomarker value and its slope,
$g(t,X_{i},\psi_{i})=\big(h(t,\psi_{i}),\frac{\partial h}{\partial t}(t,\psi_{i})\big),$
shared across transitions. Covariates $X_{i}$ are normally distributed and enter linearly through $\beta^{k^{\prime}\mid k}X_{i}$.
Longitudinal data were collected at up to $20$ random time points per subject and censored at $C_{i}\sim\mathcal{U}([10,15])$. Censoring also truncates the biomarker trajectories. The true parameter vector lies in $\mathbb{R}^{16}$. Each simulated dataset comprised $n=1000$ individuals, generating approximately balanced transition counts (Figure 4).
This trajectory mimics a biomarker whose evolution slows after treatment or onset of symptoms, while preserving continuity at the change point.
Figure 3: State transition diagram of the simulated three-state model.
6.2 Estimator Convergence and Accuracy
Parameter estimation was performed using the stochastic-gradient inference procedure of Section 4, with the Adam optimizer (learning rate $0.5$, batch size $K=100$) and an adaptive stopping criterion based on the first and second moments of parameter differences. Convergence was typically achieved in 400400500500 iterations and required around $12$ seconds per run on an AMD Ryzen 5 processor.
Across $n_{\text{runs}}=100$ independent replications, the mean relative bias of all parameters remained below $2\%$, and the root-mean-square error (RMSE) scaled approximately as $n^{-1/2}$. Parameters governing the biomarker trajectory $(\gamma,Q,R)$ were consistently well identified, while association coefficients $(\alpha^{k^{\prime}\mid k},\beta^{k^{\prime}\mid k})$ exhibited slightly higher variability for rare transitions. Representative convergence trajectories are shown in Figure 10, and quantitative results are summarized in Table 5.
@@ 6.2 Estimator Convergence and Accuracy
A short summary of the longitudinal process as well as the trajectories is given by the Figure 4 below.
@@
Figure 4: Simulated data: on the left, a sample of longitudinal measurements from $50$ individuals; on the right, observed transitions between states for the complete population of $1000$ individuals.
@@ 6.3 Estimator Convergence
To fit the model, we first have to define initial parameters, which can be particularly important for the optimization process if the model is not well-behaved. In the present case, without prior knowledge of the true parameters, we zero the values and use the identity matrix for both covariance matrices.
@@
$m_{i}^{(t)}\leftarrow\beta_{i}m_{i}^{(t-1)}+(1-\beta_{i})(\theta^{(t)}-\theta^{(t-1)})^{i},\quad\hat{m}_{i}^{(t)}=\frac{m_{i}^{(t)}}{1-\beta_{i}^{t}},$
and the optimization process is terminated when
@@
The Figure 10 shows the evolution of the parameters during the optimization process, and Table 5 summarizes the accuracy of the inferred parameters.
@@
Figure 5: Comparison of true and estimated parameters (averaged over 100 runs).
The simulation confirms that the proposed inference method accurately recovers all parameters of the multi-state joint model under realistic sample sizes. The optimization demonstrates stable convergence (Figure 10), with minimal sensitivity to initialization. The algorithm’s computational efficiency allows for extensive simulation or cross-validation studies at negligible cost.